OpenStudy (anonymous):

Someone PLEASE help me with this question! The length of a box is 5 more than the width and the height is 2 less than the width, if the volume is equal to 150 cubic ft, what are the measurements for each side?

OpenStudy (anonymous):

make the expression 1st

OpenStudy (hockeychick23):

so first make 3 equations L= w+5 H= w-2 w=w so multiply (w)(w-2)(w+5)=150 now do you know what to do?

OpenStudy (anonymous):

Yes thank you so much!!!

OpenStudy (hockeychick23):

no problem!(: